OpenStudy (anonymous):

A group of adults and students went on a class trip to Washington, DC. The number of male students was 1 more than 7 times the number of adults. The number of female students was half the number of male students. If the total number of people who went on the trip is 82, find the numbers of male students and female students.

ganeshie8 (ganeshie8):

I think you just need to setup two equations and solve

ganeshie8 (ganeshie8):

@GG1997

ganeshie8 (ganeshie8):

are you really sure the work `adult` was not a typo ?

OpenStudy (anonymous):

well how do you do that

OpenStudy (anonymous):

and I'm positive it's not a typo

ganeshie8 (ganeshie8):

idk looks i have problem interpreting the question lol

ganeshie8 (ganeshie8):

Lets start by labeling anyways let \(a\) = number of adults \(m\) = number of male students \(f\) = number of female students

ganeshie8 (ganeshie8):

`The number of male students was 1 more than 7 times the number of adults.` translates to \[m = 7a+1\tag{1}\]

ganeshie8 (ganeshie8):

`The number of female students was half the number of male students` translates to \[f=\frac{m}{2}\tag{2}\]

ganeshie8 (ganeshie8):

`the total number of people who went on the trip is 82` translates to \[a+m+f = 82\tag{3}\]

ganeshie8 (ganeshie8):

so we have a system with 3 equations, knw how to solve ?

OpenStudy (anonymous):

still struggling

OpenStudy (anonymous):

lol I really suck at math

ganeshie8 (ganeshie8):

basically we have 3 equations and we need to solve 3 unknowns : \(a, m, f\)

OpenStudy (anonymous):

ok how

ganeshie8 (ganeshie8):

\[m = 7a+1\tag{1}\] \[f=\frac{m}{2}\tag{2}\] \[a+m+f = 82\tag{3}\] rearranging second equation gives, \(m=2f\) plug this in the remaining equations

ganeshie8 (ganeshie8):

\[2f = 7a+1\tag{4}\] \[a+2f+f=82\tag{5}\] can you solve this new two equation system with two unknowns ?

OpenStudy (anonymous):

do i didvide both sides by 2

ganeshie8 (ganeshie8):

You could do that if you want to use substitution

ganeshie8 (ganeshie8):

\[2f=7a+1\] dividing 2 both sides gives \[f=\dfrac{7a+1}{2}\] you want to plug this in the other equation and solve \(a\) is it ?

ganeshie8 (ganeshie8):

that works, do it

ganeshie8 (ganeshie8):

still here ? @GG1997

OpenStudy (anonymous):

yes I'm here